Crossfire Company segments its business into two regions�East and West. The company prepared the contribution format segmented income statement shown below:

Total
Company
East West
Sales $ 900,000 $ 600,000 $ 300,000
Variable expenses 675,000 480,000 195,000
Contribution margin 225,000 120,000 105,000
Traceable fixed expenses 141,000 50,000 91,000
Segment margin 84,000 $ 70,000 $ 14,000
Common fixed expenses 59,000
Net operating income $ 25,000
Required:
1.

Compute the companywide break-even point in dollar sales. (Round intermediate calculations to 2 decimal places)

2.

Compute the break-even point in dollar sales for the East region. (Round intermediate calculations to 2 decimal places

3.

Compute the break-even point in dollar sales for the West region. (Round intermediate calculations to 2 decimal places)

4. Prepare a new segmented income statement based on the break-even dollar sales that you computed in requirements 2 and 3.
5.

Do you think that Crossfire should allocate its common fixed expenses to the East and West regions when computing the break-even points for each region?
Answer #1
Concepts and reason

Costing: It is a process of ascertaining the cost. The cost can be for a product, job, contract or service. It involves the study of cost principles, standards, and rules concerning the determination of cost of a product.

Fundamentals

Breakeven point/ Breakeven Sales Revenue: It is the point of sales where no profit and no loss are occurring. It also states that cost is equal to the income of the company. It is the point in sales dollars that can be computed by dividing fixed expenses of the company by the contribution margin ratio of the company.

The formula for Break-even Sales revenue is shown below:

BreakevenSalesRevenue=FixedexpensesContributionmarginratio{\rm{Break - even}}\,{\rm{Sales}}\,{\rm{Revenue}} = \frac{{{\rm{Fixed}}\,{\rm{expenses}}}}{{{\rm{Contribution}}\,{\rm{margin}}\,{\rm{ratio}}}}

Contribution Margin: The first profit of the company is a contribution. It is determined by subtracting the amount of variable expenses from the amount of sales. The contribution is the mandatory requirement of the organization, all indirect business expenses are dependent on this amount of contribution.

The formula of Contribution margin is shown below:

Contributionmargin=(SalesrevenueVariableexpenses){\rm{Contribution}}\,{\rm{margin}} = \left( {{\rm{Sales revenue}} - {\rm{Variable}}\,{\rm{expenses}}} \right)

Sales: It refers the revenue which is earned by the core business of the organization. The units sold during the period are multiplied by the selling price per unit in order to determine the amount of sales.

Variable expenses: These are the expenses which can be directly traceable to the product are included in the variable expenses. The units sold during the period are multiplied by the variable cost per unit in order to determine the amount of variable expenses.

Fixed Expense: The expenses which do not change with the level of activity. The expenses whose amount and nature of the occurrence, both are fixed. If the company does not produce any output, still the fixed expenses need to be paid.

(1)

Compute the Break-even point in dollar sale for the company:

Breakevenpoint=FixedcostContributionmarginratio=$200,00025%=$800,000\begin{array}{c}\\{\rm{Break - even}}\,{\rm{point}} = \frac{{{\rm{Fixed}}\,{\rm{cost}}}}{{{\rm{Contribution}}\,{\rm{margin}}\,{\rm{ratio}}}}\\\\ = \frac{{\$ 200,000}}{{25\% }}\\\\ = \$ 800,000\\\end{array}

Working note:

Compute the fixed cost:

Fixedcost=Traceablefixedexpenses+Commonfixedexpenses=$141,000+$59,000=$200,000\begin{array}{c}\\{\rm{Fixed}}\,{\rm{cost}} = {\rm{Traceable}}\,{\rm{fixed}}\,{\rm{expenses}} + {\rm{Common}}\,{\rm{fixed}}\,{\rm{expenses}}\\\\ = \$ 141,000 + \$ 59,000\\\\ = \$ 200,000\\\end{array}

Compute the company-wide contribution margin ratio:

Contributionmarginratio=ContributionmarginSales=$225,000$900,000=25%\begin{array}{c}\\{\rm{Contribution}}\,{\rm{margin}}\,{\rm{ratio}} = \frac{{{\rm{Contribution}}\,{\rm{margin}}}}{{{\rm{Sales}}}}\\\\ = \frac{{\$ 225,000}}{{\$ 900,000}}\\\\ = 25\% \\\end{array}

(2)

Compute the Break-even point in dollar sale for the region E:

Breakevenpoint=FixedcostContributionmarginratio=TraceableFixedexpensesContributionmarginratio=$50,00020%=$250,000\begin{array}{c}\\{\rm{Break - even}}\,{\rm{point}} = \frac{{{\rm{Fixed}}\,{\rm{cost}}}}{{{\rm{Contribution}}\,{\rm{margin}}\,{\rm{ratio}}}}\\\\ = \frac{{{\rm{Traceable Fixed}}\,{\rm{expenses}}}}{{{\rm{Contribution}}\,{\rm{margin}}\,{\rm{ratio}}}}\\\\ = \frac{{\$ 50,000}}{{20\% }}\\\\ = \$ 250,000\\\end{array}

Working note:

Compute the contribution margin ratio for the region E:

Contributionmarginratio=ContributionmarginSales=$120,000$600,000=20%\begin{array}{c}\\{\rm{Contribution}}\,{\rm{margin}}\,{\rm{ratio}} = \frac{{{\rm{Contribution}}\,{\rm{margin}}}}{{{\rm{Sales}}}}\\\\ = \frac{{\$ 120,000}}{{\$ 600,000}}\\\\ = 20\% \\\end{array}

(3)

Compute the Break-even point in dollar sale for the region W:

Breakevenpoint=FixedcostContributionmarginratio=TraceableFixedexpensesContributionmarginratio=$91,00035%=$260,000\begin{array}{c}\\{\rm{Break - even}}\,{\rm{point}} = \frac{{{\rm{Fixed}}\,{\rm{cost}}}}{{{\rm{Contribution}}\,{\rm{margin}}\,{\rm{ratio}}}}\\\\ = \frac{{{\rm{Traceable Fixed}}\,{\rm{expenses}}}}{{{\rm{Contribution}}\,{\rm{margin}}\,{\rm{ratio}}}}\\\\ = \frac{{\$ 91,000}}{{35\% }}\\\\ = \$ 260,000\\\end{array}

Working note:

Compute the contribution margin ratio for the region W:

Contributionmarginratio=ContributionmarginSales=$105,000$300,000=35%\begin{array}{c}\\{\rm{Contribution}}\,{\rm{margin}}\,{\rm{ratio}} = \frac{{{\rm{Contribution}}\,{\rm{margin}}}}{{{\rm{Sales}}}}\\\\ = \frac{{\$ 105,000}}{{\$ 300,000}}\\\\ = 35\% \\\end{array}

(4)

Prepare the segmented income statement of the company:

Sales
C Company
Segmented Income Statement
For the Year Ended
| Particulars Region E Region W Total Company
$250,000 $260,000

Working notes:

Compute the total sales of the company:

SalesoftheCompany=TotalsaleofregionE+TotalsaleofregionW=$250,000+$260,000=$510,000\begin{array}{c}\\{\rm{Sales}}\,{\rm{of}}\,{\rm{the}}\,{\rm{Company}} = {\rm{Total}}\,{\rm{sale}}\,{\rm{of}}\,{\rm{region}}\,{\rm{E}} + {\rm{Total}}\,{\rm{sale}}\,{\rm{of}}\,{\rm{region}}\,{\rm{W}}\\\\ = \$ 250,000 + \$ 260,000\\\\ = \$ 510,000\\\end{array}

Compute the variable expenses of region E:

VariableexpensesatBreakevensales=VariableexpensesSales×Breakevensales=$480,000$600,000×$250,000=$200,000\begin{array}{c}\\{\rm{Variable}}\,{\rm{expense}}{{\rm{s}}_{{\rm{at}}\,{\rm{Break - even}}\,{\rm{sales}}}} = \frac{{{\rm{Variable}}\,{\rm{expenses}}}}{{{\rm{Sales}}}} \times {\rm{Break - even}}\,{\rm{sales}}\\\\ = \frac{{\$ 480,000}}{{\$ 600,000}} \times \$ 250,000\\\\ = \$ 200,000\\\end{array}

Compute the variable expenses of region W:

VariableexpensesatBreakevensales=VariableexpensesSales×Breakevensales=$195,000$300,000×$260,000=$169,000\begin{array}{c}\\{\rm{Variable}}\,{\rm{expense}}{{\rm{s}}_{{\rm{at}}\,{\rm{Break - even}}\,{\rm{sales}}}} = \frac{{{\rm{Variable}}\,{\rm{expenses}}}}{{{\rm{Sales}}}} \times {\rm{Break - even}}\,{\rm{sales}}\\\\ = \frac{{\$ 195,000}}{{\$ 300,000}} \times \$ 260,000\\\\ = \$ 169,000\\\end{array}

Compute the variable expenses of the company:

VariableexpensesCompany=VariableexpensesRegionE+VariableexpensesRegionW=$200,000+$169,000=$369,000\begin{array}{c}\\{\rm{Variable}}\,\,{\rm{expense}}{{\rm{s}}_{{\rm{Company}}}} = {\rm{Variable}}\,\,{\rm{expense}}{{\rm{s}}_{{\rm{Region}}\,{\rm{E}}}} + {\rm{Variable}}\,\,{\rm{expense}}{{\rm{s}}_{{\rm{Region}}\,{\rm{W}}}}\\\\ = \$ 200,000 + \$ 169,000\\\\ = \$ 369,000\\\end{array}

Compute the contribution margin of the region E:

ContributionmarginRegionE=SalesVariableexpenses=$250,000$200,000=$50,000\begin{array}{c}\\{\rm{Contribution}}\,{\rm{margi}}{{\rm{n}}_{{\rm{Region}}\,{\rm{E}}}} = {\rm{Sales}} - {\rm{Variable}}\,\,{\rm{expenses}}\\\\ = \$ 250,000 - \$ 200,000\\\\ = \$ 50,000\\\end{array}

Compute the contribution margin of the region W:

ContributionmarginRegionW=SalesVariableexpenses=$260,000$169,000=$91,000\begin{array}{c}\\{\rm{Contribution}}\,{\rm{margi}}{{\rm{n}}_{{\rm{Region}}\,{\rm{W}}}} = {\rm{Sales}} - {\rm{Variable}}\,\,{\rm{expenses}}\\\\ = \$ 260,000 - \$ 169,000\\\\ = \$ 91,000\\\end{array}

Compute the contribution margin of the company:

ContributionmarginCompany=SalesVariableexpenses=$510,000$369,000=$141,000\begin{array}{c}\\{\rm{Contribution}}\,{\rm{margi}}{{\rm{n}}_{{\rm{Company}}}} = {\rm{Sales}} - {\rm{Variable}}\,\,{\rm{expenses}}\\\\ = \$ 510,000 - \$ 369,000\\\\ = \$ 141,000\\\end{array}

Compute the segment margin of the region E:

SegmentmarginRegionE=ContributionmarginTraceablefixedexpenses=$50,000$50,000=$0\begin{array}{c}\\{\rm{Segment}}\,{\rm{margi}}{{\rm{n}}_{{\rm{Region}}\,\,{\rm{E}}}} = {\rm{Contribution}}\,\,{\rm{margin}} - {\rm{Traceable}}\,{\rm{fixed}}\,{\rm{expenses}}\\\\ = \$ 50,000 - \$ 50,000\\\\ = \$ 0\\\end{array}

Compute the segment margin of the region W:

SegmentmarginRegionW=ContributionmarginTraceablefixedexpenses=$91,000$91,000=$0\begin{array}{c}\\{\rm{Segment}}\,{\rm{margi}}{{\rm{n}}_{{\rm{Region}}\,\,{\rm{W}}}} = {\rm{Contribution}}\,\,{\rm{margin}} - {\rm{Traceable}}\,{\rm{fixed}}\,{\rm{expenses}}\\\\ = \$ 91,000 - \$ 91,000\\\\ = \$ 0\\\end{array}

Compute the segment margin of the Company:

SegmentmarginCompany=ContributionmarginTraceablefixedexpenses=$141,000$141,000=$0\begin{array}{c}\\{\rm{Segment}}\,{\rm{margi}}{{\rm{n}}_{{\rm{Company}}}} = {\rm{Contribution}}\,\,{\rm{margin}} - {\rm{Traceable}}\,{\rm{fixed}}\,{\rm{expenses}}\\\\ = \$ 141,000 - \$ 141,000\\\\ = \$ 0\\\end{array}

Compute the net operating income or loss of the company:

NetOperatingIncome/(Loss)=SegmentmarginCommonfixedexpenses=$0$59,000=($59,000)\begin{array}{c}\\{\rm{Net}}\,{\rm{Operating}}\,{\rm{Income/}}\left( {{\rm{Loss}}} \right) = {\rm{Segment}}\,{\rm{margin}} - {\rm{Common}}\,{\rm{fixed}}\,{\rm{expenses}}\\\\ = \$ 0 - \$ 59,000\\\\ = \left( {\$ 59,000} \right)\\\end{array}

While computing the break-even point of each region, company C should not allocate the cost of common fixed expenses between the E region and W region.

Ans: Part 1

The break-even point in sales dollar for the company is $800,000.

Part 2

The break-even point in sales dollar for the region E is $250,000.

Part 3

The break-even point in sales dollar for the region W is $260,000.

Part 4

The net operating loss of the company is $59,000.